In a local hardware store the price of the store brand indoor house paint is $18 per gallon. The daily volume of sales is 30 gallons. The price elasticity of demand for the paint is estimated to be -3. Determine the demand equation for the paint assuming that it is linear.
Answer #1

P = $ 18 and Q = 30 gallons

Q = a - bP (linear demand curve)

dQ/dP x P/Q = - 3 (price elasticity)

- b x 18/30 = -3

b = 5

30 = a - 5 x 18

a = 120

Demand curve: Q = 120 - 5P